MONT 100N – Modeling the Environment Second Essay

The second of the larger paper assignments this semester will deal with various aspects of the Purgatory Chasm state park that we will visit on October 21.

For the purposes of this assignment, you should think of yourself as something like an author of a guidebook entry for Purgatory Chasm, writing for educated readers who know nothing about the history of the chasm or the possible scientific explanations of how its geological features came to be, but who might be interested in visiting and want to know what’s interesting about this location. For the assignment, you will need to conduct some research and write an essay incorporating answers to the questions posed below. If you need help, our Science Librarian, Ms. Barbara Merolli will be more than happy to assist you in locating sources in print and online forms.

As a responsible guide book author, you should try to indicate why the chasm is an interesting and enjoyable place to visit, but you shouldn’t go overboard with outlandish claims such as, “this is the most amazing place on the face of the earth.” Keep the tone quietly enthusiastic, but not overblown. Most importantly, your paper must include answers to the specific questions below, but it should be structured as an actual essay, not just a list of answers. Document your sources with footnotes and a “Works Consulted” list at the end of your paper.

Note: The assignment will be perfectly doable even if you are not able to participate in the field trip. However, the personal experience gained from actually seeing the place will probably be helpful. If you go on the trip, you will probably want to take pictures to refresh your memory as you write later. It is OK to share photos with members of the class who cannot participate in the field trip. But of course, the essay itself must be your own work.

Your finished essay is due no later than 5:00pm on Friday, November 3, by email to

jlittle@holycross.edu.

Questions

1. Who previously owned the land purchased by the state of Massachusetts to form the current state park?

2. When was the Purgatory Chasm State Reservation created and how was it created?

3. How big is the state park?

4. What are the names of the six trails and five of the marked rock formations? (You might include descriptions and/or personal impressions of which are the most picturesque here.)

5. Who was Thomas Clemens Wohlbrück and what was his connection with Purgatory Chasm?

6. What theory of the formation of the Chasm was proposed by Edward Hitchcock in 1841? What is the problem with Hitchcock’s theory?

7. What theory for the formation of the Chasm was proposed by Prof. W.O. Crosby and his son, Irving B. Crosby. What is the problem with Prof. Crosby’s theory?

8. What theory for the formation of the Chasm was proposed by Prof. R.J. Lougee? What is the problem with Prof. Lougee’s theory?

9. What theory for the formation of the Chasm was proposed by Prof. Mauri Pelto?

Guidelines for this assignment

Follow these specifications:

1. You will need to consult sources – books, articles, web sites, etc. for this assignment. If you make use of direct quotations or ideas presented in those sources, then provide enough information for an interested reader to find and check the source you are using. For books, give the author, title, publisher, year of publication, and page where your information is coming from. For web sites, give the full URL, an author if you can identify one, and the date you consulted the site.

2. The article (paper) as a whole should be well-organized and “flow” from one point to the next. For instance, you might start out with a few sentences about the general history of Purgatory Chasm, then describe the rock formations and trails that are the attractions for visitors. Then your final section might deal with the different scientific theories of how the Chasm was formed. Do not get ahead of yourself. Say one thing at a time. If you decide you want to add something, do not just tack it on somewhere. Instead, figure out where it fits in the overall scheme of your description and put it where it belongs.

3. Your sentences should be clear, readable, and not overly wordy. You should aim for a lively style and the writing should be in your own voice. But please think twice before using very colloquial ways of saying things. If they contribute to the point you are trying to make, that is fine. But do it for a reason. Also, please do not use a thesaurus to find impressive-sounding words when simpler alternatives are available. An important lesson to learn for writing in college is that thesauri do not capture the nuances in meanings that underlie good writing. If you must use a thesaurus, also look up any words you select in a dictionary to make sure they mean exactly what you think they do!

4. Errors in mechanics (spelling, punctuation, grammar) and usage should not annoy the reader (your editor). (A word of warning: Your reader (i.e. Prof. Little) tends to get easily annoyed by this sort of thing!) Try as much as possible to avoid those mistakes.